The Letters of Miss Jerusha Abbott

Mr. Daddy-Long-Legs Smith 215 FERGUSSEN HALL 24th September

Dear Kind-Trustee-Who-Sends-Orphans-to-College,

Here I am! I travelled yesterday for four hours in a train. It’s a funny sensation, isn’t it? I never rode in one before.

College is the biggest, most bewildering place–I get lost whenever I leave my room. I will write you a description later when I’m feeling less muddled; also I will tell you about my lessons. Classes don’t begin until Monday morning, and this is Saturday night. But I wanted to write a letter first just to get acquainted.

It seems queer to be writing letters to somebody you don’t know. It seems queer for me to be writing letters at all–I’ve never written more than three or four in my life, so please overlook it if these are not a model kind.

Before leaving yesterday morning, Mrs. Lippett and I had a very serious talk. She told me how to behave all the rest of my life, and especially how to behave towards the kind gentleman who is doing so much for me. I must take care to be Very Respectful.

But how can one be very respectful to a person who wishes to be called John Smith? Why couldn’t you have picked out a name with a little personality? I might as well write letters to Dear Hitching-Post or Dear Clothes-Prop.

I have been thinking about you a great deal this summer; having somebody take an interest in me after all these years makes me feel as though I had found a sort of family. It seems as though I belonged to somebody now, and it’s a very comfortable sensation. I must say, however, that when I think about you, my imagination has very little to work upon. There are just three things that I know:

I. You are tall. II. You are rich. III. You hate girls.

I suppose I might call you Dear Mr. Girl-Hater. Only that’s rather insulting to me. Or Dear Mr. Rich-Man, but that’s insulting to you, as though money were the only important thing about you. Besides, being rich is such a very external quality. Maybe you won’t stay rich all your life; lots of very clever men get smashed up in Wall Street. But at least you will stay tall all your life! So I’ve decided to call you Dear Daddy-Long-Legs. I hope you won’t mind. It’s just a private pet name we won’t tell Mrs. Lippett.

The ten o’clock bell is going to ring in two minutes. Our day is divided into sections by bells. We eat and sleep and study by bells. It’s very enlivening; I feel like a fire horse all of the time. There it goes! Lights out. Good night.

Observe with what precision I obey rules–due to my training in the John Grier Home.

Yours most respectfully, Jerusha Abbott